### v1

# NAMO-LLM: Efficient Navigation Among Movable Obstacles with Large Language Model Guidance

Abstract
Several planners have been proposed to compute robot paths that reach desired goal regions while avoiding obstacles. However, these methods fail when all pathways to the goal are blocked. In such cases, the robot must reason about how to reconfigure the environment to access task-relevant regions—a problem known as Navigation Among Movable Objects (NAMO). While various solutions to this problem have been developed, they often struggle to scale to highly cluttered environments. To address this, we propose NAMO-LLM, a sampling-based planner that searches over robot and obstacle configurations to compute feasible plans specifying which obstacles to move, where, and in what order. Its key novelty is a non-uniform sampling strategy guided by Large Language Models (LLMs) biasing the tree construction toward directions more likely to yield a solution. We show that NAMO-LLM is probabilistically complete and demonstrate through experiments that it efficiently scales to cluttered environments, outperforming related works in both runtime and plan quality.
I Introduction
Several planning algorithms have been developed to compute robot paths from an initial configuration to a desired one while avoiding obstacles [1, 2, 3, 4, 5]. However, in highly cluttered environments, obstacle-free paths to goal regions may not exist. In such cases, rather than reporting task failure, a more effective approach is to enable robots to interact with their environment by repositioning movable obstacles. This requires planning algorithms capable of determining which obstacles to move, in what order, and where to reposition them to create a feasible path to the goal; see Fig. 1. This planning problem, known as Navigation Among Movable Objects (NAMO) [6], has been proven to be NP-hard [7].
Several algorithms have been proposed to address NAMO problems. Early approaches relied on heuristic solutions which lacked completeness guarantees [8]. To overcome this, search-based planning methods have been introduced. These methods either provide completeness guarantees for specific subclasses of NAMO problems or sacrifice computational efficiency to ensure completeness across a broader range of scenarios [6, 9, 10, 11]. For instance, [6] addressed NAMO problems where disconnected free-space components can be connected by moving a single obstacle at a time. This formulation allows the problem to be represented as a graph structure, enabling the use of resolution-complete planners. This work was later extended to handle cases where at most $k$ obstacles must be moved, with each obstacle requiring only a single displacement [10]. A more general approach is proposed in [11] to solve non-monotone NAMO problems, where multiple obstacles may need to be repositioned—potentially more than once—to establish connectivity between disconnected free-space components. While this algorithm is complete under mild assumptions, its computational cost scales exponentially with the number of movable objects. Extensions of these approaches to account for physical interactions between objects are presented in [12, 13], while extensions to unknown environments are considered in [14, 15, 16, 17, 18, 19]. In addition to search-based methods, sampling-based planners have been developed to improve efficiency while still maintaining (probabilistic) completeness guarantees for general NAMO problems [20, 21, 22, 23]. The key idea in these approaches is to incrementally construct trees by sampling and adding nodes, where each node represents the environment’s configuration, and edges model feasible manipulation actions. However, their efficiency declines as the number of movable objects or the number of required obstacle relocations grows, due to their uniform sampling strategy.
Figure 1: A ground robot is tasked with reaching a destination behind the building block; see (1). All direct pathways are initially blocked (see Fig. 2). The robot autonomously removes obstacles to clear a path; see (2)–(4).
This paper addresses this limitation by designing a new sampling-based NAMO algorithm that efficiently scales to environments characterized by large numbers of movable obstacles and/or the need for multiple obstacle relocations to reach desired regions. Inspired by [22], our algorithm incrementally builds trees that explore the space of obstacle and robot configurations. Each tree node represents an environment configuration, while edges among nodes correspond to manipulation actions that relocate movable obstacles. At each iteration, a tree node and a manipulation action are sampled. Application of that action to the sampled tree node results in new tree node that is added to the tree if feasible. The key novelty of our work lies in a non-uniform sampling strategy that leverages Large Language Models (LLMs) to accelerate plan generation. Specifically, our sampling strategy directs the tree to grow towards LLM-recommended directions at certain times while allowing random exploration otherwise. We refer to the resulting planner as NAMO-LLM, for Navigation Among Movable Obstacles with Large Language Model guidance. We show that NAMO-LLM is probabilistically complete meaning that the probability of computing a feasible plan converges to one as the number of iterations goes to infinity. We conduct extensive comparative experiments against both search-based planners [18] and sampling-based planners relying on uniform sampling strategies [22], demonstrating that our method significantly reduces both the time required to compute the first feasible plan and the corresponding number of obstacles that need to be relocated. We also demonstrate NAMO-LLM on hardware experiments; see Fig. 1.
Related Works: In addition to the NAMO algorithms discussed earlier, our work is also related to research on rearrangement planning (RP) problems where robots manipulate movable obstacles [24, 25, 26, 27]. RP differs from NAMO because it considers predefined goal positions for the movable obstacles and typically does not specify a final goal for the robot. While our work primarily focuses on NAMO problems, we believe that our non-uniform sampling strategy can also be integrated with various sampling-based approaches for planning among movable obstacles, including RP and manipulation planning [28, 29]. However, a formal integration with these approaches is beyond the scope of this paper. Furthermore, several works design non-uniform sampling strategies to accelerate sampling-based planners [30, 31, 32, 33, 34, 35, 36]. For example, [34] learns a non-uniform sampling distribution from demonstrations. However, we emphasize that these works target different planning problems from ours, focusing on designing obstacle-free paths between an initial and a desired robot configuration (while also assuming that such paths exist). As such, they cannot be directly applied to our setting. It is worth noting that, unlike [32, 33, 34, 35, 36], our work does not require training neural networks; instead, ours leverages pre-trained LLMs.
Summary of Contributions: First, we introduce NAMO-LLM, a novel sampling-based planner designed to efficiently solve NAMO problems in highly cluttered environments. Second, to the best of our knowledge, this is the first integration of LLMs within a sampling-based NAMO planner, significantly accelerating plan generation. Third, we conduct extensive experiments demonstrating that our approach outperforms existing planning algorithms in terms of efficiency and solution quality. Fourth, we release an open-source implementation of our proposed algorithm [37].
II Problem Formulation
Robot Modeling: We consider a robot with the following kinematics: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ where $\mathbf{p}(t)\in{\mathcal{P}}$ and $\mathbf{u}(t)\in{\mathcal{U}}$ denote the robot configuration (e.g., position and orientation) and the control input at time step $t$, respectively. We assume that ${\mathbf{p}}(t)$ is known at all times and that the robot is holonomic and, therefore, it can move along any desired path. The robot is also equipped with a gripper, enabling it to manipulate one movable object at a time.
Figure 2: Top-view representation of the urban environment shown in Fig. 1. The set ${\mathcal{P}}$ is defined by the position of the center of the disk-shaped robot. Given ${\mathcal{N}}(0)$, the set ${\mathcal{O}}_{M}(0)$ includes only obstacles 1-6.
Environment Modeling: The robot operates in an environment $\Omega$ containing a set of known obstacles ${\mathcal{O}}=\{o_{1},\dots,o_{O}\}$. The obstacles are classified as movable obstacles, collected in a set ${\mathcal{O}}_{\text{movable}}\subseteq{\mathcal{O}}$, and non-movable ones. A movable object can be repositioned by the robot as long as it first grasps it. The configuration (e.g., position and orientation) of an obstacle $o_{i}\in{\mathcal{O}}$ at time $t$ is denoted by $c_{i}(t)\in{\mathcal{C}}_{i}$ where ${\mathcal{C}}_{i}$ denotes its configuration space. We also denote by $co_{i}(t)$ the region in the configuration space $\mathcal{P}$ of the robot that corresponds to configurations where the robot is in collision with obstacle $o_{i}$; see Fig. 2. Thus, the robot’s free space at time $t$, denoted by ${\mathcal{P}}_{\text{free}}(t)$, is defined as ${\mathcal{P}}_{\text{free}}(t)={\mathcal{P}}\setminus\mathcal{CO}(t)$ where $\mathcal{CO}(t)=\cup_{i=1}^{n_{O}}\{co_{i}(t)\}$. Notice that ${\mathcal{P}}_{\text{free}}(t)$ may consist of multiple connected components whose boundaries are determined by the configuration of the obstacles. We denote by ${\mathcal{N}}(t)\subseteq{\mathcal{P}}_{\text{free}}(t)$, the free-space component that ${\mathbf{p}}(t)$ lies in; see Fig. 2.
Robot Task: The robot is tasked with reaching a desired configuration starting from an initial one ${\mathbf{p}}(0)$ while avoiding collisions with obstacles. Hereafter, we denote by ${\mathcal{P}}_{g}\subseteq{\mathcal{P}}$ a set collecting all desired robot configurations. We consider cases where all configurations in ${\mathcal{P}}_{g}$ are initially unreachable due to the initial configuration of obstacles, i.e., ${\mathcal{P}}_{g}\cap{\mathcal{N}}(0)=\emptyset$. In such cases, the robot must determine which movable obstacles need to be relocated and where they should be placed. To formalize this, we define the set of manipulable obstacles at time $t$ as $\mathcal{O}_{M}(t)\subseteq{\mathcal{O}}_{\text{movable}}$ which includes all movable obstacles $o_{i}$ that the robot can directly reach, i.e., $\mathcal{O}_{M}(t)=\{o_{i}\in{\mathcal{O}}_{\text{movable}}\mid co_{i}(t)\cap{\mathcal{N}}(t)\neq\emptyset\}$; see also Fig. 2. Next, we define manipulation actions of the form $a(c_{i}(t),c_{i}^{\prime})$ which involve moving a manipulable object $o_{i}\in\mathcal{O}_{M}(t)$, from its current configuration $c_{i}(t)$ to a new configuration $c_{i}(t^{\prime})=c_{i}^{\prime}$, for some $t^{\prime}>t$. In other words, the application of this action requires the robot to approach the obstacle’s current location (determined by $c_{i}(t)$), grasp it, and relocate it to the new configuration $c_{i}^{\prime}$. Thus, our goal is to compute a plan $\tau$, i.e., a sequence of manipulation actions so that eventually ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t)\neq\emptyset$. We defined this plan as $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, where $\tau(k)$ refers to the $k$-th manipulation action, $k\in\{0,1,\dots,H\}$, for some horizon $H\geq 0$. Note that the index $k$ is different from the time step $t$, as the application of an action $\tau(k)$ may require multiple time steps $t$. The problem that this paper addresses can be summarized as follows:
Problem 1
Given an initial robot configuration ${\mathbf{p}}(0)$, a set ${\mathcal{P}}_{g}$ of desired configurations, and initial obstacle configurations $c_{i}(0)$, compute a manipulation plan $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, so that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$.
III Proposed Planning Algorithm
In this section, we present NAMO-LLM, our proposed planner to address Problem 1. In Section III-A, we present a sampling-based planner to design plans $\tau$. In Section III-B, we present a non-uniform sampling strategy that accelerates the plan synthesis phase by leveraging Large Language Models (LLM). In Section III-C, we discuss the input prompt to the LLM guiding the sampling strategy. In Section III-D, we show that NAMO-LLM is probabilistically complete.
Input: (i) initial robot configuration ${\mathbf{p}}(0)$; (ii) initial obstacle configuration ${\mathbf{c}}(0)$; (iii) Set ${\mathcal{O}}_{\text{movable}}$ ; (iv) goal set ${\mathcal{X}}_{g}$
Output: The manipulation plan $\tau$
1 Compute ${\mathcal{N}}(0)$;
2 ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ ;
3 Initialize ${\mathcal{V}}=\{{\mathbf{q}}_{0}\},~{}{\mathcal{E}}=\emptyset,~{}{\mathcal{X}}_{g}=\emptyset,~{}m=1$ ;
4 while $m<m_{\text{max}}$ do
5       Sample a node ${\mathbf{q}}_{\text{rand}}\in{\mathcal{V}}$ using $f_{\mathcal{V}}$;
6       $[{\mathbf{q}}_{\text{new}},a(c_{i}(t),c_{i}^{\prime})]=\texttt{TreeExpansion}({\mathbf{q}}_{\text{rand}})$;
7       if  ${\mathbf{q}}_{\text{new}}\neq\varnothing$ then
8             ${\mathcal{V}}\leftarrow{\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$, ${\mathcal{E}}\leftarrow{\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$;
9             $A(({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}}))=a(c_{i}(t),c_{i}^{\prime})$;
10             Extract ${\mathcal{N}}(t^{\prime})$ from ${\mathbf{q}}_{\text{new}}=[\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$;
11             if  ${\mathcal{P}}_{g}\cap\mathcal{N}(t^{\prime})\neq\emptyset$  then
12                   ${\mathcal{X}}_{g}\leftarrow{\mathcal{X}}_{g}\cup\{{\mathbf{q}}_{\text{new}}\}$;
13
14      Update iteration index: $m=m+1$;
15
16if ${\mathcal{X}}_{g}\neq\emptyset$ then
17       Pick a node ${\mathbf{q}}_{g}\in{\mathcal{X}}_{g}$;
18       Compute path $\tau_{{\mathbf{q}}}$ over ${\mathcal{G}}$ from ${\mathbf{q}}(0)$ to ${\mathbf{q}}_{g}$;
19       Using $A$ and $\tau_{{\mathbf{q}}}$, compute the manipulation plan $\tau$ ;
20
21else
22       No plan $\tau$ has been found;
23
Algorithm 1 NAMO-LLM
1 Compute the set of manipulable obstacles ${\mathcal{O}}_{M,\text{rand}}(t)$ ;
2 Sample a manipulable obstacle $o_{i}$ from a given distribution $f_{{\mathcal{O}}_{M}}$ ;
3 Sample a random configuration $c_{i}^{\prime}$ from $f_{{\mathcal{C}}_{i}}$;
4 if $co_{i}^{\prime}\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ and $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$  then
5       if $c_{i}(t)\to c^{\prime}_{i}$ is collision-free  then
6             Initialize ${\mathbf{c}}(t^{\prime})={\mathbf{c}}(t)$ and update $c_{i}(t^{\prime})=c_{i}^{\prime}$ in ${\mathbf{c}}(t^{\prime})$;
7             $\mathcal{P}_{\text{free}}(t^{\prime})=\mathcal{P}_{\text{free}}(t)\cup co_{i}(t)\setminus co_{i}^{\prime}$ ;
8             Given ${\mathbf{c}}(t^{\prime})$, select $\mathcal{N}(t^{\prime})$ ;
9             ${\mathbf{q}}_{\text{new}}=(\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime}))$ ;
10             return ${\mathbf{q}}_{\text{new}}$, $a(c_{i}(t),c_{i}^{\prime})$
11      else
12             return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
13else
14       return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
Algorithm 2 TreeExpansion(${\mathbf{q}}_{\text{rand}}$)
III-A Sampling-based Planning
We propose a sampling-based planner to address Problem 1 summarized in Alg. 1 The proposed algorithm incrementally builds a tree that explores the configuration space of the movable objects and the robot’s free space components to compute a manipulation plan $\tau$ as defined in Section II. We denote the constructed tree by ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where ${\mathcal{V}}$ and ${\mathcal{E}}$ is a set of nodes and edges, respectively. The set ${\mathcal{V}}$ contains states of the form ${\mathbf{q}}(t)=[\mathcal{N}(t),{\mathbf{c}}(t)]$, where ${\mathbf{c}}(t)$ is a vector stacking the configuration of all obstacles, i.e., ${\mathbf{c}}(t)=[c_{1}(t),c_{2}(t),\dots,c_{n}(t)]$ and ${\mathcal{N}}(t)$ denotes the free-space component that ${\mathbf{p}}(t)$ lies in. The root of the tree models the initial configuration of the obstacles and the robot, i.e., ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ [lines 1-1, Alg. 1]. The tree is initialized so that ${\mathcal{V}}=\{{\mathbf{q}}(0)\}$ and ${\mathcal{E}}=\emptyset$ [line 1, Alg. 1]. Also, the tree is built incrementally by sampling new states ${\mathbf{q}}_{\text{new}}$ that are added to the tree structure [line 1, Alg. 1]. The algorithm terminates after a user-specified number of iterations $m_{\text{max}}$ and returns a feasible plan $\tau$ (if it has been found). In what follows, we present a detailed description of our method.
At every iteration $m$ of Alg. 1, we perform the following two steps. First, we sample from a given discrete mass function $f_{{\mathcal{V}}}:{\mathcal{V}}\rightarrow[0,1]$ a tree node denoted by ${\mathbf{q}}_{\text{rand}}(t)=[{\mathcal{N}}_{\text{rand}}(t),{\mathbf{c}}_{\text{rand}}(t)]\in{\mathcal{V}}$ [line 1, Alg. 1]. The mass function $f_{{\mathcal{V}}}$ defines the probability of selecting a node ${\mathbf{q}}\in{\mathcal{V}}$. Second, we sample a new node ${\mathbf{q}}_{\text{new}}$ that will be added to the tree; see Alg. 2. Specifically, given ${\mathbf{q}}_{\text{rand}}$, we compute the corresponding set of manipulable obstacles denoted by $\mathcal{O}_{M,\text{rand}}(t)$, [line 2, Alg. 2]. Then, we sample from a discrete mass function $f_{{\mathcal{O}}_{M}}:{\mathcal{O}}_{M,\text{rand}}\rightarrow[0,1]$ a manipulable obstacle $o_{i}\in\mathcal{O}_{M}(t)$ [line 2, Alg. 2]. The mass function $f_{{\mathcal{O}}_{M}}$ defines the probability of selecting a manipulable obstacle $o_{i}\in{\mathcal{O}}_{M,\text{rand}}$. Given the selected manipulable obstacle $o_{i}$, we sample a new configuration $c_{i}^{\prime}\in{\mathcal{C}}_{i}$ using a density function $f_{\mathcal{C}_{i}}:\mathcal{C}_{i}\rightarrow[0,\infty)$ [line 2, Alg. 2]. Designing the functions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ is deferred to Section III-B.
Figure 3: Illustration of the free-space components ${\mathcal{N}}(t)$ (left column) and ${\mathcal{N}}^{1}(t)$ (right column) in two environments.
Next, we check if a collision-free path exists for the robot to move $o_{i}$ from $c_{i}(t)$ to $c_{i}^{\prime}$ [line 2, Alg. 2]. Two necessary conditions must be met for $c_{i}^{\prime}$ to admit such a collision-free path; if they are not satisfied then such a path does not exist [line 2, Alg. 2]. First, $c_{i}^{\prime}$ must not intersect other obstacles, i.e., $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}^{\text{rand}}(t)\setminus co_{i}(t)$, where ${\mathcal{P}}_{\text{free}}^{\text{rand}}(t)$ is computed using ${\mathbf{q}}_{\text{rand}}$. Second, $co_{i}^{\prime}$ must not lie in a component of ${\mathcal{P}}_{\text{free}}$ that becomes accessible only if an obstacle $o_{j}\neq o_{i}$ is relocated first. To formalize this, we define ${\mathcal{P}}^{i}_{\text{free}}(t)$ as the robot’s free space if obstacle $o_{i}$ is removed from the environment, i.e., ${\mathcal{P}}^{i}_{\text{free}}(t)={\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$. Correspondingly, ${\mathcal{N}}^{i}(t)$ denotes the component of ${\mathcal{P}}^{i}_{\text{free}}(t)$ containing the robot’s position ${\mathbf{p}}(t)$; see Fig. 3. Thus, the second condition requires $co_{i}^{\prime}$ to belong to ${\mathcal{N}}^{i}_{\text{rand}}(t)$ (i.e., to ${\mathcal{N}}^{i}(t)$ computed using ${\mathbf{q}}_{\text{rand}}$).
If the above two conditions hold, then there may exist such a collision-free path for $o_{i}$. This path is defined as a sequence of obstacle configurations, i.e., $\boldsymbol{\zeta}=c_{i}(t),c_{i}(t+1),\dots,c_{i}(t+s),\dots,c_{i}(t^{\prime})$ such that (i) $c_{i}(t^{\prime})=c_{i}^{\prime}$ and (ii) $co_{i}(t+s)\in{\mathcal{N}}^{i}(t)$, for all $s\in\{1,2,\dots,t^{\prime}-t\}$. This can be checked using any existing motion planner such as RRT or A* [3]. If such a path $\boldsymbol{\zeta}$ exists, this means the obstacle $o_{i}$ can be moved to $c_{i}^{\prime}$ giving rise to ${\mathbf{c}}(t^{\prime})$ that differs from ${\mathbf{c}}(t)$ only in the configuration of the obstacle $o_{i}$ [line 2, Alg. 2]. In general, ${\mathcal{N}}^{i}(t)$ may split into multiple components once $o_{i}$ is moved to $c_{i}^{\prime}$, each adjacent to $co_{i}^{\prime}$. The robot can be in any of these components, and we randomly select one to be ${\mathcal{N}}(t^{\prime})$ [line 2, Alg. 2].
Given the selected free-space component ${\mathcal{N}}(t^{\prime})$ and ${\mathbf{c}}(t^{\prime})$, we construct the node ${\mathbf{q}}_{\text{new}}=[{\mathcal{N}}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$ [line 2, Alg. 2] and the sets of nodes and edges are updated as ${\mathcal{V}}={\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$ and ${\mathcal{E}}={\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$, respectively [lines 1, Alg. 1]. Also, once an edge $({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ is added to the tree, we store the corresponding manipulation $a(c_{i}(t),c_{i}^{\prime})$ that can enable it. We formalize this by iteratively constructing a function $A$ which returns the corresponding action for each edge in ${\mathcal{E}}$ [line 1, Alg. 1]. If a collision-free path $\boldsymbol{\zeta}$ does not exist or if $c_{i}^{\prime}$ is not a valid configuration, then we set $q_{\text{new}}=\varnothing$ and the iteration $m$ will terminate without adding any new nodes to the tree [lines 2-2, Alg. 2].
Once a node ${\mathbf{q}}_{\text{new}}$ is added to the tree, we check if ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t^{\prime})\neq\emptyset$. If that holds, it means that the robot can reach its goal configuration set ${\mathcal{P}}_{g}$. The tree nodes that satisfy this property are collected in a set ${\mathcal{X}}_{g}\subseteq{\mathcal{V}}$ [line 1, Alg. 1] that is initialized as ${\mathcal{X}}_{g}=\emptyset$ [line 1, Alg. 1]. Once the algorithm terminates, we select a tree node ${\mathbf{q}}_{g}\in{\mathcal{X}}_{g}$ (if it is non-empty) [lines 1-1, Alg. 1]. Then, we compute the path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(H+1)$ over ${\mathcal{G}}$ connecting ${\mathbf{q}}_{g}$ to the tree root ${\mathbf{q}}(0)$, i.e., $\tau_{{\mathbf{q}}}(0)={\mathbf{q}}(0)$, $\tau_{{\mathbf{q}}}(H+1)={\mathbf{q}}_{g}$, and $\tau_{{\mathbf{q}}}(k)\in{\mathcal{V}}$, $\forall k\in\{1,\dots,H+1\}$ [line 1, Alg. 1]. Using $\tau_{{\mathbf{q}}}$ and the function $A$, we can extract the plan $\tau=\tau(0),\dots,\tau(H)$ modeling the sequence of manipulation actions that need to be applied to solve Problem 1 [line 1, Alg. 1]. The node ${\mathbf{q}}_{g}$ can be selected randomly or based on any user-defined criterion such as the horizon $H$ of $\tau$ or the distance that needs to be traveled to execute $\tau$; see also Rem. III.1.
Remark III.1 (Execution of $\tau$)
Each action $\tau(k)=a(c_{i},c_{i}^{\prime})$ in $\tau$ involves moving an object $o_{i}$ from its current configuration $c_{i}$ to a new configuration $c_{i}^{\prime}$. Execution of this action should be in accordance with the corresponding path $\tau_{{\mathbf{q}}}$. This means that the robot should grasp $o_{i}$, while being in ${\mathcal{N}}$ as determined by $\tau_{{\mathbf{q}}}(k)=[{\mathcal{N}},{\mathbf{c}}]$, and drop-it off at $c_{i}^{\prime}$ while being in ${\mathcal{N}}^{\prime}$ determined by $\tau_{{\mathbf{q}}}(k+1)=[{\mathcal{N}}^{\prime},{\mathbf{c}}^{\prime}]$. According to line 2 in Alg. 2, there must exist a collision-free path for the robot to execute this movement. Also, observe that, as in [22], our planner decouples the computation of the robot’s motion from the computation of obstacle movements. Specifically, it does not determine the exact grasp poses for the objects of interest but instead tracks the free-space component of the robot.
Figure 4: Input prompt for GPT-4o and $K=1$ in Case Study V (Sec. IV).
III-B LLM-guided Sampling Strategy
The sampling-based planner above relies on the distributions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ to generate ${\mathbf{q}}_{\text{new}}$. While these can be uniform, as e.g., in [22], such strategies often lead to slow plan synthesis, particularly as environmental complexity (e.g., the number of movable objects) increases. To accelerate the planning process, we propose non-uniform mass functions $f_{{\mathcal{V}}}$ and $f_{{\mathcal{O}}_{M}}$ that bias sampling toward directions where a feasible solution is more likely to be found. Our approach uses LLMs to identify these promising directions.
First, we define $f_{{\mathcal{V}}}$ as follows:
$f_{{\mathcal{V}}}({\mathbf{q}}\mid{\mathcal{V}},{\mathcal{V}}_{v})=\begin{cases}\frac{p_{\text{rand}}}{|{\mathcal{V}}_{v}|}+\frac{1-p_{\text{rand}}}{|{\mathcal{V}}|},&\text{if}~{}{\mathbf{q}}\in{\mathcal{V}}_{v}\\ \frac{1-p_{\text{rand}}}{|{\mathcal{V}}|},&\text{otherwise}\end{cases}$ (1)
where $p_{\text{rand}}\in[0,1)$ is a user-specified parameter and ${\mathcal{V}}_{v}\subseteq{\mathcal{V}}$; ${\mathcal{V}}_{v}$ will be defined later once $f_{{\mathcal{O}}_{M}}$ is defined. According to (1), a node ${\mathbf{q}}\in{\mathcal{V}}_{v}$ is selected to be the node ${\mathbf{q}}_{\text{rand}}$ with probability $p_{\text{rand}}$ while any node ${\mathbf{q}}\in{\mathcal{V}}$ is selected with probability $1-p_{\text{rand}}$. Notice that if $p_{\text{rand}}=0$, then $f_{{\mathcal{V}}}$ models uniform sampling. As $p_{\text{rand}}$ increases, the sampling strategy becomes more biased toward selecting nodes ${\mathbf{q}}\in{\mathcal{V}}_{v}$.
Second, we define the function $f_{{\mathcal{O}}_{M}}$ as follows:
$f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})=\begin{cases}p_{\text{obs}}+\frac{1}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}=o_{\text{LLM}},\\ \frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}\in\mathcal{O}_{M,\text{rand}},\end{cases}$ (2)
where $p_{\text{obs}}\in[0,1)$ is a user-specified parameter and $o_{\text{LLM}}$ is manipulable obstacle, i.e., $o_{\text{LLM}}\in{\mathcal{O}}_{M}$, that is recommended by an LLM to relocate. A discussion on how to prompt an LLM to select $o_{\text{LLM}}$ is deferred to Section III-C. According to (2), the obstacle $o_{\text{LLM}}\in{\mathcal{O}}_{M}$ is selected for relocation with probability $p_{\text{obs}}$ while any obstacle $o_{i}\in{\mathcal{O}}_{M}$ is selected with probability $1-p_{\text{obs}}$. The probability of selecting $o_{i}\notin{\mathcal{O}}_{M}$ is 00. Observe that if $p_{\text{obs}}=0$, then $f_{{\mathcal{O}}_{M}}$ models uniform sampling. As $p_{\text{obs}}$ increases, the sampling strategy becomes more biased toward selecting $o_{\text{LLM}}$. Also, we note that it is possible that the LLM may fail to select a manipulable obstacle or it may generate nonsensical information. In such cases, we set $p_{\text{obs}}=0$ for that iteration so that a manipulable obstacle is selected randomly; see Section III-C.
Next, we define the set ${\mathcal{V}}_{v}$ used for the construction of $f_{{\mathcal{V}}}$ in (1). First, we define a function $v:{\mathcal{V}}\rightarrow\mathbb{N}$ assigning a positive score to each node ${\mathbf{q}}\in{\mathcal{V}}$. This function is initialized as $v({\mathbf{q}}(0))=0$. Second, we show how the score of a newly added node ${\mathbf{q}}_{\text{new}}$ is generated. If ${\mathbf{q}}_{\text{new}}$ was constructed by relocating $o_{\text{LLM}}$, then we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})+1$. Otherwise, we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})$. Informally, $v({\mathbf{q}}_{\text{new}})$ counts how many nodes along the path connecting ${\mathbf{q}}_{\text{new}}$ to ${\mathbf{q}}(0)$ have been constructed based on LLM recommendations. Third, we construct the set ${\mathcal{V}}_{v}$ so that it collects the nodes with the highest score $v$, i.e., ${\mathcal{V}}_{v}=\{{\mathbf{q}}\in{\mathcal{V}}~{}|~{}v({\mathbf{q}})=\max_{{\mathbf{q}}^{\prime}\in{\mathcal{V}}}v({\mathbf{q}}^{\prime})\}$.
Finally, we select a uniform density function for $f_{{\mathcal{C}}_{i}}$. Non-uniform density functions $f_{{\mathcal{C}}_{i}}$ can also be implemented similarly, e.g., by querying an LLM to select configurations $c_{i}^{\prime}$ for which collision-free paths exist; for simplicity, we abstain from this design choice.
III-C Prompt Structure
In this section, we describe how we query a pre-trained LLM to select a manipulable obstacle $o_{\text{LLM}}$. The input prompt to the LLM consists of three parts (see Fig.4): (1) Task description which provides information about the environment’s structure (i.e., the location of all movable and non-movable obstacles and the goal region) and specifies that the robot’s task is to remove obstacles so that the goal region is reachable; (2) Rules that the robot must follow to accomplish its task. Specifically, these rules state that the robot can only remove manipulable obstacles, should minimize the number of obstacles removed and should avoid picking manipulable obstacles that are blocked by other obstacles (see e.g., Case Study VI in Section IV); (3) Output format which defines how the LLM’s response should be formatted for compatibility with Alg. 2. In particular, the LLM is required to generate $K\geq 1$ obstacles in the form $\text{obs}_{i}$ where $i$ is a unique identifier of each obstacle. Among the generated obstacles, one is selected randomly and denoted by $o_{\text{LLM}}$ in (2). Note that the LLM may make content and format errors. A content error occurs when the selected obstacle is not manipulable, and a format error occurs when the LLM’s output does not comply with the required format. If either of these errors occurs, we set $p_{\text{obs}}=0$.
III-D Probabilistic Completeness
In what follows, we show that NAMO-LLM is probabilistically complete. To formally state this result, we first introduce the notion of path clearance [22]. Consider any path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(k),\dots,\tau_{{\mathbf{q}}}(H+1)$ computed by Alg. 1 yielding a plan $\tau$ solving Problem 1. We say that $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$ if any alternative path $\tilde{\tau}_{{\mathbf{q}}}=\tilde{\tau}_{{\mathbf{q}}}(0),\dots,\tilde{\tau}_{{\mathbf{q}}}(k),\dots,\tilde{\tau}_{{\mathbf{q}}}(H+1)$ satisfying the following conditions (i)-(ii) for all $k\in\{1,\dots,H\}$ also corresponds to a feasible plan $\tilde{\tau}$. (a) The transitions from $\tau_{{\mathbf{q}}}(k-1)$ to $\tau_{{\mathbf{q}}}(k)$ and from $\tilde{\tau}_{{\mathbf{q}}}(k-1)$ to $\tilde{\tau}_{{\mathbf{q}}}(k)$ must involve relocating the same manipulable obstacle $o_{i}$, albeit potentially to different configurations. (b) Let $c_{i}^{k}$ and $\tilde{c}_{i}^{k}$ denote the configurations of that obstacle $o_{i}$ in $\tau_{{\mathbf{q}}}(k)$ and $\tilde{\tau}_{{\mathbf{q}}}(k)$, respectively. These configurations must satisfy $\lVert c_{i}^{k}-\tilde{c}_{i}^{k}\rVert\leq\epsilon$ for all $k$, for some $\epsilon>0$. In what follows, we refer to the satisfaction of conditions (i)–(ii) concisely as $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$.
Theorem III.2
Assume that there exists a feasible path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),...,\tau_{{\mathbf{q}}}(H+1)$ with clearance $\epsilon>0$. If $p_{\text{rand}}\in[0,1)$, $p_{\text{obs}}\in[0,1)$, and the motion planner used to check existence of collision-free paths between obstacle configurations in line 2 of Alg. 2 is complete, then the probability that NAMO-LLM will find a feasible path approaches $1$ as $m$ goes to $\infty$.
Proof:
To show this result, we need first to introduce the following notations and definitions. Let ${\mathcal{G}}_{m}=\{{\mathcal{V}}_{m},{\mathcal{E}}_{m}\}$ denote the tree constructed by Algorithm 1 at the end of iteration $m$. Since the plan $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$, there exist alternative feasible plans $\tilde{\tau}_{{\mathbf{q}}}$ such that $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$ for all $k$. Throughout the proof, any reference to a state $\tilde{\tau}_{{\mathbf{q}}}(k)$ assumes this $\epsilon$-closeness to the plan $\tau_{{\mathbf{q}}}$.
We also define the following two random variables that will be used to model the sampling process and the computation of a feasible plan. First, we define Bernoulli random variables $X_{m}$ that are equal to $1$ if the following holds for any state $\tilde{\tau}_{{\mathbf{q}}}(k)$ and $k\in\{0,1,\dots,H+1\}$ at iteration $m$: (i) a state $\tilde{\tau}_{{\mathbf{q}}}(k)$ is added to the tree (i.e., $\tilde{\tau}_{{\mathbf{q}}}(k)\in{\mathcal{V}}_{m}$) via an edge from a predecessor $\tilde{\tau}_{{\mathbf{q}}}(k-1)\in{\mathcal{V}}_{m^{\prime}}$ for some $m^{\prime}<m$ (i.e., $(\tilde{\tau}_{{\mathbf{q}}}(k-1),\tilde{\tau}_{{\mathbf{q}}}(k))\in{\mathcal{E}}_{m}$), and (ii) no state $\tilde{\tau}_{{\mathbf{q}}}^{\prime}(k)$ satisfying $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}^{\prime}(k)\rVert\leq\epsilon$ was already present in the tree prior to iteration $m$ (i.e., $\tilde{\tau}_{{\mathbf{q}}}^{\prime}(k)\notin{\mathcal{V}}_{m^{\prime}}$ for all $m^{\prime}<m$). Additionally, once a state $\tilde{\tau}_{{\mathbf{q}}}(H+1)$ satisfying both (i) and (ii) is added at an iteration $m$, no future iterations can satisfy condition (ii) for any $k$. In this case, by convention, we set $X_{m^{\prime}}=1$, for all $m^{\prime}>m$ (i.e., the probability of success of $X_{m^{\prime}}$ is $1$).
Second, we define the random variable $Y_{m}$ counting the number of successes of the Bernoulli variables $X_{1},X_{2},\dots,X_{m}$, i.e., $Y_{m}=\sum_{i=1}^{m}X_{i}$. By construction of $X_{i}$, if $Y_{m}>H$, then a state $\tilde{\tau}_{{\mathbf{q}}}(H+1)$ has been added to ${\mathcal{V}}_{m}$ indicating that a solution has been found. Therefore, to show probabilistic completeness, it suffices to show that the probability $\mathbb{P}(Y_{m}\geq H+1)$ approaches $1$ as $m\to\infty$.
Using the Chebyshev inequality, we can bound the “failure” probability, i.e., the probability that fewer than $H$ successful expansions occurred in $m$ iterations (which means the goal has not been reached), as follows:
$\mathbb{P}(Y_{m}\leq H)$ $\leq\mathbb{P}\Bigl{(}|Y_{m}-\mathrm{E}(Y_{m})|\geq\mathrm{E}(Y_{m})-H\Bigl{)}$
$\leq\frac{\mathrm{Var}(Y_{m})}{(\mathrm{E}(Y_{m})-H)^{2}}$ (3)
The expectation $\mathrm{E}(Y_{m})$ and variance $\mathrm{Var}(Y_{m})$ of $Y_{m}$ are defined as follows
$\mathrm{E}(Y_{m})=\sum_{i=1}^{m}\mathrm{E}(X_{i})=\sum_{i=1}^{m}p_{i}$ (4)
and
$\mathrm{Var}(Y_{m})=\sum_{i=1}^{m}\mathrm{Var}(X_{i})+2\sum_{1\leq i<j\leq m}\mathrm{Cov}(X_{i},X_{j})$
$=\sum_{i=1}^{m}p_{i}(1-p_{i})+2\sum_{1\leq i<j\leq m}\Bigl{[}\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\Bigr{]},$ (5)
where $p_{i}$ is the probability of success of $X_{i}$.
In what follows, we show (a) $\lim_{m\to\infty}E(Y_{m})=\infty$ and (b) $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$. Using (b), we can re-write (III-D) as follows:
$\mathbb{P}(Y_{m}\leq H)\leq\frac{\mathrm{Var}(Y_{m})}{(\mathrm{E}(Y_{m})-H)^{2}}\leq\frac{3\mathrm{E}(Y_{m})}{(\mathrm{E}(Y_{m})-H)^{2}}.$ (6)
Then combining (a) with (6), we get that $\lim_{m\to\infty}\mathbb{P}(Y_{m}\leq H)=0$. This equivalently means $\lim_{m\to\infty}\mathbb{P}(Y_{m}\geq H+1)=1$ completing the proof.
Proof of (a): Here we show that $\lim_{m\to\infty}E(Y_{m})=\infty$. Step 1: To show this, we first show the probability of success of $X_{m}$, denoted by $p_{m}$, is strictly positive, i.e., $p_{m}>0$, for all $m$. This result will later be used to establish a lower bound on $\mathbb{P}(Y_{m}\geq H+1)$. Showing this consists of the following two parts.
Part 1: First, let $\hat{{\mathcal{V}}}_{m}\subseteq{\mathcal{V}}_{m}$ collect all states $\tilde{\tau}_{{\mathbf{q}}}(k)\in{\mathcal{V}}_{m}$ for which there does not exist a state $\tilde{\tau}_{{\mathbf{q}}}(k^{\prime})$ with $k^{\prime}>k$ that is already in the set ${\mathcal{V}}_{m}$. A necessary requirement for $X_{m}=1$ is to select a state from $\hat{{\mathcal{V}}}_{m-1}$ to be ${\mathbf{q}}_{\text{rand}}$. By construction of $f_{{\mathcal{V}}}({\mathbf{q}}\mid{\mathcal{V}},{\mathcal{V}}_{v})$, the probability $p_{1}^{m}=f_{{\mathcal{V}}}({\mathbf{q}}\mid{\mathcal{V}},{\mathcal{V}}_{v})$ of selecting a state ${\mathbf{q}}_{\text{rand}}\in\hat{{\mathcal{V}}}_{m-1}$ is strictly positive. Specifically, we have that $p_{1}^{m}\geq\min\{\frac{p_{\text{rand}}}{|{\mathcal{V}}_{v}|}+\frac{1-p_{\text{rand}}}{|{\mathcal{V}}_{m-1}|},\frac{1-p_{\text{rand}}}{|{\mathcal{V}}_{m-1}|}\}=\frac{1-p_{\text{rand}}}{|{\mathcal{V}}_{m-1}|}\geq\frac{1-p_{\text{rand}}}{m}$ where the last inequality is due to $|{\mathcal{V}}_{m-1}|\leq{m}$. Observe that $p_{2}^{m}>0$ since by assumption $p_{\text{rand}}<1$.
Part 2: Second, assume that a node $\tilde{\tau}_{{\mathbf{q}}}(k)$ from $\hat{{\mathcal{V}}}_{m}$ has been selected to be ${\mathbf{q}}_{\text{rand}}$. Under this assumption, there exist nodes ${\mathbf{q}}_{\text{new}}$ that can be added to the tree satisfying conditions (i)-(ii) required for $X_{m}=1$. Specifically, all nodes $\tilde{\tau}_{{\mathbf{q}}}(k+1)$ are such candidate nodes. Next, we show that the probability $p_{2}^{m}$ of sampling a state $\tilde{\tau}_{{\mathbf{q}}}(k+1)$ to be $q_{\text{new}}$ is strictly positive, i.e., $p_{2}^{m}>0$, for all $m$. Given ${\mathbf{q}}_{\text{rand}}=\tilde{\tau}_{{\mathbf{q}}}(k)$, sampling a state ${\mathbf{q}}_{\text{new}}=\tilde{\tau}_{{\mathbf{q}}}(k+1)$ will occur under the following two requirements (as per the $\epsilon$-clearance definition). First, at state ${\mathbf{q}}_{\text{rand}}=\tilde{\tau}_{{\mathbf{q}}}(k)$ the obstacle $o_{i}$ that will be selected to relocate is the one selected in the transition from $\tau_{{\mathbf{q}}}(k)$ to $\tau_{{\mathbf{q}}}(k+1)$. By construction of $f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})$, since the obstacle $o_{i}$ is manipulable (by construction of $\tau_{{\mathbf{q}}}$) this will occur with probability $p_{o_{i}}^{m}\geq\min\{p_{\text{obs}}+\frac{1}{|\mathcal{O}_{M,\text{rand}}|},\frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|}\}$. Observe that since $p_{\text{obs}}<1$, we have that $p_{o_{i}}^{m}>0$ for all $m$. Second, assuming that $o_{i}$ will be selected for relocation, its new configuration, denoted by $\tilde{c}_{i}^{k+1}$, should satisfy $\lVert\tilde{c}_{i}^{k+1}-c_{i}^{k+1}\rVert{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq\epsilon}$ where $c_{i}^{k+1}$ is the configuration of $o_{i}$ at state $\tau_{{\mathbf{q}}}(k+1)$. Since by assumption $\epsilon>0$ and $f_{{\mathcal{C}}_{i}}$ is a uniform density function, we get that the probability $p_{c_{i}}^{m}$ of sampling a $\tilde{c}_{i}^{k+1}$ that satisfies this requirement is also strictly positive for all $m$. Thus, we get that $p_{2}^{m}=p_{o_{i}}^{m}p_{c_{i}}^{m}>0$. Consequently, there must exist a positive constant $p_{2}^{\text{min}}$ such that $p_{2}^{m}\geq p_{2}^{\text{min}}$ for all $m\geq 0$ We note also since a complete planner is used to check if there exists a collision-free path reaching $\tilde{c}_{i}^{k+1}$ from $\tilde{c}_{i}^{k}$ (i.e., the configuration of the obstacle $o_{i}$ in $\tilde{\tau}_{{\mathbf{q}}}(k)$), then the sampled configuration $\tilde{c}_{i}^{k+1}$ will be never be rejected in line 2 incorrectly (i.e., if such a collision-free path exists).
Combining Part 1 and 2, we get that the probability $p_{m}$ of success of $X_{m}$ is strictly positive since $p_{m}=p_{1}^{m}p_{2}^{m}\geq\frac{(1-p_{\text{rand}})p_{2}^{m}}{m}\geq\frac{(1-p_{\text{rand}})p_{2}^{\text{min}}}{m}>0.$
Step 2: Given the result from Step 1, we can re-write the expected value of $Y_{m}$ in (4) as:
$E(Y_{m})=\sum_{i=1}^{m}p_{1}^{i}p_{2}^{i}\geq\sum_{i=1}^{m}\frac{(1-p_{\text{rand}})p_{2}^{\text{min}}}{i}=\alpha\sum_{i=1}^{m}\frac{1}{i}$
where $\alpha=(1-p_{\text{rand}})p_{2}^{\text{min}}>0$ is a positive constant across all iterations $m$. Then,
$\mathrm{E}(Y_{m})\geq\alpha\sum_{i=1}^{m}\frac{1}{i}$
When $m\to\infty$, $\sum_{i=1}^{m}\frac{1}{i}\to\infty$. Therefore, we get $\mathrm{E}(Y_{m})\to\infty$ completing the proof of part (a).
Proof of (b): Here we show that $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$. Recall from (III-D) that the variance of $Y_{m}$ is given by:
$\mathrm{Var}(Y_{m})$ $=\sum_{i=1}^{m}p_{i}(1-p_{i})+2\sum_{1\leq i<j\leq m}\Bigl{[}E(X_{i}X_{j})-p_{i}p_{j}\Bigr{]}$ (7)
First, we focus on the term $\sum_{i=1}^{m}p_{i}(1-p_{i})$ in (7). Since $0<p_{i}\leq 1$, we have that
$\sum_{i=1}^{m}p_{i}(1-p_{i})<\sum_{i=1}^{m}p_{i}=\mathrm{E}(Y_{m})$
Second, we focus on the term $\sum_{1\leq i<j\leq m}\Bigl{[}E(X_{i}X_{j})-p_{i}p_{j}\Bigr{]}$ in (7). Since $X_{i}X_{j}\leq X_{i}$, we have that:
$\mathrm{E}(X_{i}X_{j})\leq\mathrm{E}(X_{i})=p_{i}$
Hence,
$\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\leq p_{i}(1-p_{j})<p_{i}$
Then,
$2\sum_{1\leq i<j\leq m}\Bigl{[}\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\Bigr{]}<2\sum_{i=1}^{m}p_{i}=2\mathrm{E}(Y_{k})$
Thus we get $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$ completing the proof of part (b). ∎
IV Experiments
In this section, we present extensive numerical experiments to evaluate the efficiency of our method. Specifically, we demonstrate that our approach can effectively solve NAMO problems involving hundreds of movable obstacles and requiring the manipulation of multiple obstacles, significantly outperforming recent search-based methods [18] and sampling-based planners relying on uniform sampling [22] in terms of both runtime and plan quality. Hardware demonstrations of NAMO-LLM on ground robots are provided in [38]. For these experiments, we use the GPT-4o and Gemini-1.5-flash models. All experiments were conducted on a computer with an Intel Core i5 2.4GHz processor and 8 GB of RAM. Software implementation of NAMO-LLM can be found in [37].
IV-A Setting Up Comparative Experiments
NAMO Baselines: We evaluate NAMO-LLM against two baselines: (i) RandomTree, a sampling-based approach from [22] that uses uniform sampling strategies, and (ii) NAMO-SA, a search-based method from [18]. Notably, RandomTree can be recovered from our method by setting $p_{\text{rand}}=p_{\text{obs}}=0$. The approach in [18] selects obstacles for removal by computing the total distance from the robot to the obstacle and from the obstacle to the goal region using the $\text{A}^{*}$ algorithm. This is done for all manipulable objects, and the obstacle with the lowest total distance is selected. The chosen obstacle is then moved to a pre-defined storage room if a collision-free path towards that storage room exists. If no path to the storage room exists, the manipulable obstacle with the next lowest cost is selected; more details can be found in [18]. In our implementation of [18], we model the storage room as an external region adjacent to a specified portion of the workspace boundary. Once an obstacle reaches this boundary segment, it is assumed to be instantaneously transferred to the storage room. On the other hand, our method designs custom drop-off locations for the obstacles during the sampling process. Note that this favors the baseline as our method may sample invalid drop-off locations (e.g., configurations lying in the obstacle space), resulting in sample rejection; see lines 2-2 in Alg. 2. Also in [18], once an obstacle is moved to the storage room, it is never considered for relocation again. This design choice favors that baseline, as the number of obstacles it needs to handle decreases over time—unlike our method. Thus, the maximum horizon $H$ of a plan designed by [18] is equal to the number $n$ of movable obstacles, while in ours it is possible that $H>n$ since our planner is allowed to reconfigure obstacles multiple times. We also note here that [18] includes re-planning strategies to handle unknown obstacles (i.e., obstacles with unknown locations and movability). Since our focus is on known environments, we exclude these capabilities. Finally, to ensure fair comparisons, our implementation of NAMO-LLM ([line 2, Alg. 2]), [18], and [22] share the same RRT planner to check if there exists a collision-free path towards the drop-off location of an obstacle and the storage area.
Robot and Obstacle Configuration Space: We consider a disk-shaped robot with radius $r=0.2$ (Case Studies I-IV) and $r=0.5$ (Case Studies V-VI) and configuration space ${\mathcal{P}}$ determined by the position of the center of the robot. We consider polygonal obstacles with configuration space ${\mathcal{C}}_{i}$ defined by their position and orientation.
Evaluation Metrics: As evaluation metrics, we use (i) the runtime to compute the first feasible plan and (ii) the horizon $H$ of the designed plan. Runtimes and plan horizons $H$ for our method and the sampling-based baseline [22] are averaged over $100$ trials.
[…]
IV-B Comparative Experiments against NAMO Baselines
In what follows, we present six case studies that differ in their complexity in terms of the total number $n$ of movable obstacles, the initial configuration of the obstacles, and the minimum horizon, denoted by $H_{\text{min}}$, of a plan $\tau$ required to reach ${\mathcal{P}}_{g}$. In our method, NAMO-LLM$(p_{\text{rand}},p_{\text{obs}})$, we employ distinct parameter settings for $p_{\text{rand}}$ and $p_{\text{obs}}$. The following results are obtained using the pre-trained GPT 4o model.
Method $|{\mathcal{V}}|$ Time (secs) $H$
NAMO-SA [18] N/A 99.05 27
RandomTree [22] 15.3 14.94 4.3
K=1 NAMO-LLM(0.2, 0.2) 6.6 6.42 3.1
NAMO-LLM(0.2, 0.8) 3.15 3.21 2.5
NAMO-LLM(0.8, 0.2) 6.26 5.74 3.74
NAMO-LLM(0.8, 0.8) 2.71 2.88 2.52
K=2 NAMO-LLM(0.2, 0.2) 7.85 8.77 3.6
NAMO-LLM(0.2, 0.8) 3.30 4.91 2.65
NAMO-LLM(0.8, 0.2) 5.7 4.94 3.5
NAMO-LLM(0.8, 0.8) 3.11 4.46 2.89
K=3 NAMO-LLM(0.2, 0.2) 7.05 7.68 3.10
NAMO-LLM(0.2, 0.8) 3.30 4.47 2.43
NAMO-LLM(0.8, 0.2) 5.45 5.14 3.35
NAMO-LLM(0.8, 0.8) 3.21 4.25 2.67
[…]
Method $|{\mathcal{V}}|$ Time (secs) $H$
[…]
RandomTree [22] 34.90 42.78 4.57
5
NAMO-LLM(0.2, 0.8) 3.86 9.99 2.75
NAMO-LLM(0.8, 0.8) 4.47 7.80 3.04
[…]
RandomTree [22] 92.72 71.32 5.72
5
NAMO-LLM(0.2, 0.8) 6.33 8.52 3.70
NAMO-LLM(0.8, 0.8) 4.37 5.91 3.74
[…]
RandomTree [22] 290.83 408.68 7.41
5
NAMO-LLM(0.2, 0.8) 8.23 19.31 4.16
NAMO-LLM(0.8, 0.8) 5.70 17.12 4.5
[…]
Case Study I ($n=50,H_{\text{min}}=1$): In this case study, we consider a $50\times 50$ environment with $n=50$ movable obstacles. The initial configurations of the obstacles and the robot, as well as the desired robot configuration, are depicted in Fig. 5. The environment is designed such that a plan with horizon $H_{\text{min}}=1$ exists to reach ${\mathcal{P}}_{g}$, meaning that the minimum number of obstacles that must be reconfigured to reach ${\mathcal{P}}_{g}$ is $1$. We test NAMO-LLM for various values of the parameters $p_{\text{rand}}$, $p_{\text{obs}}$, and $K$. The comparative results are summarized in Table I. Observe that our method significantly outperforms both baselines in terms of both the runtime required to compute the first feasible plan and the horizon $H$ of that plan. For instance, the runtimes of the search- and sampling-based methods are $99.05$ and $14.94$ seconds, respectively, while our method can compute the first feasible plan significantly faster; e.g., the average runtime for $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$ is $2.88$ seconds. The average horizon of the plan computed by our method is consistently less than $4$, which is comparable to that of the sampling-based baseline, while the horizon of the search-based one is $27$. We note that the performance of our method does not change significantly with respect to $K\in\{1,2,3\}$. We attribute this to the fact that there are multiple candidate obstacles near ${\mathcal{P}}_{g}$ that can be removed to get access to it and that the employed LLM is capable of identifying them.
Case Study II ($n=100,H_{\text{min}}=1$): In this case study, we consider a $65\times 65$ environment containing $n=100$ movable obstacles, where ${\mathcal{P}}_{g}$ can be reached by removing at least one obstacle; see Fig 5. The comparative results are shown in Table II. Due to space limitations, we present results for our method only with $K=1$. Observe that the performance gap between our method and the baselines increases as $n$ grows. For instance, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, our method finds the first feasible plan in $7.80$ seconds with an average horizon of $H=3.04$. In contrast, [22] takes $42.78$ seconds to compute a plan with a slightly longer horizon ($H=4.57$), while [18] requires $876.36$ seconds and produces a plan with $H=65$.
Case Study III ($n=50,H_{\text{min}}=2$): Here we consider a $50\times 50$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5. The results in Table II show that the performance gap between our method and the baselines further increases as the minimum required horizon $H_{\text{min}}$ grows.
Case Study IV ($n=100,H_{\text{min}}=2$): This case study examines a more complex environment with dimensions $65\times 65$ and $n=100$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5. The results are summarized in Table II, showing that the performance gap further increases as both $n$ and the minimum horizon $H_{\text{min}}$ increase. For instance, our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, finds the first feasible plan in $17.12$ seconds with an average horizon of $H=4.5$. In contrast, the sampling-based approach takes $408.68$ seconds to compute a plan with a longer horizon of $7.41$, while the search-based method requires $877.97$ seconds and produces a plan with $H=65$.
[…]
Case Study V ($n=50,H_{\text{min}}=6$): In this case study, we consider a $40\times 40$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least six obstacles (see Fig. 6). This scenario is significantly more complex due to the larger number of obstacles that need to be relocated. This complexity is also reflected in the increased runtimes reported in Table III, compared to the previous case studies. Our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, computed the first feasible plan in $48.22$ seconds, with an average horizon of $H=26.48$. In contrast, [22] failed to find a feasible plan within $6$ hours while [18] can identify a feasible plan within $141.91$ seconds with $H=42$.
Case Study VI ($n=10,H_{\text{min}}=3$): Finally, we consider a more complex scenario to illustrate the utility of parameter $K$, where—unlike previous case studies—some manipulable obstacles are blocked by others and cannot be relocated until the blocking ones are moved. Specifically, we consider a $20\times 20$ environment with $n=10$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires relocating the manipulable obstacle $o_{1}$. However, unlike Case Studies I-IV, there do not exist collision-free paths to any possible configuration $c_{1}^{\prime}$ unless both $o_{8}$ and $o_{2}$ are moved first; see Fig 6. Thus, in this case, at least $3$ obstacles must be relocated to reach ${\mathcal{P}}_{g}$. The comparative results are shown in Table IV. Unlike Case Study I, the performance of our method improves significantly for $K\in\{2,3\}$ compared to $K=1$. This occurs because, with $K=1$, the LLM often prioritizes removing $o_{1}$ first (i.e., $o_{\text{LLM}}=o_{1}$ in (2)), even though it cannot be moved due to being blocked by $o_{8}$ and $o_{2}$. Consequently, any configuration $c_{1}^{\prime}$ sampled for $o_{1}$ is rejected by line 2 in Alg. 2, slowing down the overall algorithm. However, when $K>1$, the LLM generates a set of obstacles that need to be removed to access ${\mathcal{P}}_{g}$, which may include $o_{2}$ and/or $o_{8}$. Randomly selecting an obstacle from this set helps the algorithm avoid getting stuck due to persistent incorrect LLM recommendations and, consequently, improves performance. Also, observe that our method outperforms [22], which relies on random exploration, in terms of runtime—even when configured with $K=1$. We attribute this to the fact that the LLM occasionally provides correct guidance (i.e., recommending the removal of $o_{2}$ and $o_{8}$ before $o_{1}$) after some redundant actions have been applied (e.g., relocating $o_{3}$ first), as these actions influence the input prompt to the LLM. Overall, our method outperforms both baselines when $K$ is $2$ or $3$. Observe in Table IV that [18] eventually moved all obstacles to the storage area leading to plans with $H=10$.
Summary of Comparisons: Overall, the performance gap between NAMO-LLM and the baselines increases as $n$ or $H_{\text{min}}$ grows. In general, we observed that our method tends to achieve its lowest runtime for computing the first feasible plan when configured with $p_{\text{rand}}=p_{\text{obs}}=0.8$ and $K=1$; if manipulable obstacles are not blocked by other obstacles; see e.g., Case Studies I–V. We attribute this to the LLM’s ability to correctly reason about which obstacle to remove in these scenarios. However, when $p_{\text{rand}}=0.2$ and $p_{\text{obs}}=0.8$, the generated plans tend to have a shorter horizon $H$. This occurs because a smaller $p_{\text{rand}}$ enables the algorithm to select tree nodes for expansion more uniformly, leading to better-quality first plans, albeit with slightly higher runtimes. We also observed that pre-trained LLMs may occasionally make infeasible recommendations regarding which manipulable obstacle to relocate, especially when these obstacles are blocked by others and cannot be repositioned until the blocking ones are moved first; see e.g., Case Study VI. In such cases, setting $K>1$ is beneficial, as it prevents the algorithm from getting stuck due to poor LLM recommendations. Finally, we note that these results depend on the selected LLM.
Method $|{\mathcal{V}}|$ Time (secs) $H$
NAMO-SA [18] N/A 141.91 42
RandomTree [22] over 30k over 22k N/A
NAMO-LLM(0.2, 0.8) 106.12 143.83 19.65
NAMO-LLM(0.8, 0.8) 34.78 48.22 26.48
[…]
Method $|{\mathcal{V}}|$ Time (secs) $H$
NAMO-SA [18] N/A 87.28 10
RandomTree [22] 326.65 227.46 8.95
K=1 NAMO-LLM(0.2, 0.2) 112.45 188.95 10.1
NAMO-LLM(0.2, 0.8) 32.8 359.60 8.11
NAMO-LLM(0.8, 0.2) 85.6 162.09 10.8
NAMO-LLM(0.8, 0.8) 18.45 187.31 8.54
K=2 NAMO-LLM(0.2, 0.2) 82.9 81.92 10.6
NAMO-LLM(0.2, 0.8) 18.5 52.66 7.95
NAMO-LLM(0.8, 0.2) 32.3 37.69 11.1
NAMO-LLM(0.8, 0.8) 12.7 31.05 9.5
K=3 NAMO-LLM(0.2, 0.2) 70.95 98.02 10.65
NAMO-LLM(0.2, 0.8) 26.25 93.51 8.35
NAMO-LLM(0.8, 0.2) 34.9 44.23 12.05
NAMO-LLM(0.8, 0.8) 12.65 42.47 9.9
[…]
Method $|{\mathcal{V}}|$ Time (secs) $H$
[…]
RandomTree [22] 290.83 408.68 7.41
NAMO-LLM(0.2, 0.8) 84.35 233.94 15.8
NAMO-LLM(0.8, 0.8) 24.75 104.30 17.2
GPT 4o 5.70 17.12 4.16
[…]
RandomTree [22] over 30k over 22k N/A
NAMO-LLM(0.2, 0.8) 27.70 21.03 9.6
NAMO-LLM(0.8, 0.8) 15.15 14.24 12.15
GPT 4o 34.78 48.22 19.65
TABLE V: Effect of LLM Model on average tree size, runtimes, and horizon. The line ‘GPT 4o’ reports the best results of our method when paired with GPT 4o.
IV-C Comparative Results with an Alternative LLM
In this section, we demonstrate the performance of our method when paired with the Gemini-1.5-flash model on Case Studies IV ($n=100$, $H_{\text{min}}=2$) and V ($n=50$, $H_{\text{min}}=6$). The comparative results are shown in Table V. In Case Study IV, our method outperforms both baselines in terms of runtime but produces longer plans than the sampling-based baseline (while still shorter than those of the search-based one). In Case Study V, however, our approach achieves better performance than both baselines in terms of both runtime and plan horizon $H$. Overall, we observe that switching from GPT-4o to Gemini-1.5-flash tends to improve the performance of our algorithm in scenarios where manipulable obstacles are blocked by other obstacles (see Case Study V) and deteriorate its performance in cases with large number of movable obstacles (see Case Study IV). Nevertheless, our method still remains more scalable than the baselines.
V Conclusion
In this paper, we proposed NAMO-LLM, a probabilistically complete sampling-based planner for NAMO problems. Through comparative experiments, we demonstrated its scalability and efficiency in highly cluttered environments. Future work will focus on extending NAMO-LLM to settings with obstacles of unknown location and movability.